# EPFL-Smart-Kitchen-30: Densely annotated cooking dataset with 3D kinematics to challenge video and language models

### v1

@@ 1 Introduction
Table 1: Action dataset comparison. ##\## indicates "number of" for simplicity. The remaining columns mark following features: parametric model for motion representation (PM), structured actions (SA), markerless video recording (ML), depth recording (DR), and absolute positioning (AP).
@@ 3.2 Data collection procedure
To capture realistic cooking scenarios, subjects prepared a meal from reading a recipe to cleaning up, leading to significantly longer recordings than most existing action datasets (Table 1). We recruited 16 subjects (four males and twelve females, two left-handed, ages 20-46) to cook for up to five sessions in the EPFL-Smart-Kitchen (see Supp. Mat.). During each session, subjects are asked to follow one of four different recipes (omelet, pad thai, risotto, ratatouille), adapted to their preferences and requirements. Overall, we recorded and processed 29.7h of cooking experiments, corresponding to 320,760 frames per camera for 49 cooking sessions. All procedures were approved by the EPFL-Ethical Board. Subjects’ faces are anonymized across all videos to address privacy concerns. All subjects consented and were informed about the ethics (see Supp. Mat.).
@@ 4.2 Action recognition benchmark
Results. Simple concatenation of all modalities as inputs yielded slightly better performance compared to single video modality (Table 4.2). Meanwhile, transfer learning from a ViT model trained on EPIC-KITCHENS-100 (20) boosted the performance for baselines with visual inputs and reduced the performance for pose only. Furthermore, adding (hand and body) pose information to the video-based models yielded better overall performance, which mainly benefits from the verb prediction improvements. However, simply concatenating tokens from different modalities barely improved the performance. To efficiently utilize multiple modalities without significant computational cost, we integrate egocentric view, multi-exocentric
@@ 4.4 Situated full-body motion generation benchmark
Table 4: Full-body motion generation results. Models trained and evaluated on the EPFL-Smart-Kitchen-30. FID:Fréchet Inception Distance, DIV:Diversity, MM:Multimodality, MMd:Multimodal distance.
Pre-processing, Baselines and Metrics To achieve robust full-body motion representation, we combine joint locations and angles for the body, hands, and eye gaze, resulting in a 327-dimensional redundant motion representation. We process fine-grained action text with linguistic tags and extract egocentric visual features with CLIP’s text and image encoders (92). As baselines, we adapt two strong motion generation models, T2M-GPT (15) and MoMask (16), training them with verb-noun pairs and verb-only prompts (see Supp. Mat.). Like in HumanML3D (26), we train the quantitative evaluator on EPFL-Smart-Kitchen-30 to measure the R-precision top-1 to top-3 (T1 to T3), multimodal distance (MMd), Fréchet Inception Distance (FID), Diversity (DIV), and Multimodality (MM).
Results. Our qualitative analysis revealed compelling examples of successfully generated full body motion sequences that accurately represent natural cooking behaviors (Figure 5 and Supp. Mat.). Quantitatively, we found that just like on HumanML3D(26), MoMask (16) consistently outperforms T2M-GPT (15) in different settings and metrics (Table 4.4). Furthermore, when conditioning on the egocentric view, the model was able to find some clues to minimize the global distribution and thus make the FID score lower.
@@ 5 Conclusion, future work and impact
Overall, we share multi-view, multimodal action understanding, modeling, and video question answering benchmarks to leverage the potential of multi-modality for improving action understanding and to fuel foundation models. This is particularly interesting for emerging multi-modal models (85, 86, 93).
@@ A.1 Details on devices and sensors
We installed nine Kinect Azure cameras (67) inside the kitchen and asked the subjects to wear the HoloLens 2 headset (68) to record video streams. The RGB streams were recorded at 30 FPS, and the depth streams were recorded at 10 FPS. Meanwhile, we also extracted the head position, hand poses, and eye gaze data at 30 FPS from the HoloLens 2. Apart from the video streams, we also installed six wired IMU sensors on some of the appliances (fridge/cupboards) and two wireless IMU sensors on the frequently used tools (knife/spatula). Those IMU sensors were recorded at 100 FPS (Figure A.1 B).
@@ A.2 Synchronization and calibration of the sensors
As is shown in Figure A.1 C, we synchronized the egocentric and exocentric cameras using the offset of a flashing light at the beginning of the session and use the recorded timestamps to match the captured frames. The Kinect Azure cameras can output regular audio signals; we used those signals to synchronize all the Kinect Azure cameras and the IMU sensors. We validated the calibration and synchronization by projecting the regressed 3D body poses to all the views and manually checking if the poses were aligned with the body in all the views at randomly selected timestamps.
@@ A.3 List of recipes
Fridge Counter 2$\times$×Stoves Ventilation Sink Drawers
@@ F.1.5 Question engineering
Q6: I am currently [VERB] the [NOUN], what [BE||||past,future] [NEXT/PREVIOUS] action ?
@@
Q7: I am currently [VERB], what [BE||||past,future] [NEXT/PREVIOUS] action ?
@@
Q8: I am currently [ACTIVITY], what [BE||||past,future] [NEXT/PREVIOUS] activity?
@@
Q11: I am currently [VERB] the [NOUN], what [BE||||past,present,future] my [NEXT/CURRENT/PREVIOUS] activity ?
@@
Q12: I am currently [VERB], what [BE||||past,present,future] my [NEXT/CURRENT/PREVIOUS] activity ?
@@ F.4 Details on the Action Recognition Benchmark
Data from different modalities are first sent to different modality-specific projection layers to obtain the query tokens and then sent to the transformer backbone to extract cross-modality token-wise features. Finally, features from different modalities are merged using Average Pooling to make action predictions. Specifically, for video-type data, we follow the original implementation of VideoMAE (8), select 16 frames and split into 16 $\times$ 16 patches for each frame; we introduce the egocentric view and one global exocentric view as inputs in our baseline designs. For point-type data (e.g., hand-body poses and eye gaze), we sample poses of 32 frames. For every pose, we concatenate all the pose joints and their corresponding confidences together and apply a linear projection layer to generate a query token. We tried two different variants that either use VideoMAE (8) weights pre-trained on the EPIC-KITCHENS-100 (20) or trained from scratch. We follow standard practice (48, 98) to predict verb class and noun class separately and then apply the outer product to obtain the action class.
@@ F.4.2 Data Preprocessing
We adopt VideoMAE (8) model with the VIT-L backbone (99) to build our multimodal baselines. Data from different modalities are projected into query tokens, processed by a transformer for feature extraction, and merged via Average Pooling for action prediction. Video data uses VideoMAE (8) with 16 frames, incorporating egocentric and global exocentric views. Point-type data (e.g., hand-body poses, eye gaze) samples 32-frame poses, concatenates joint positions with confidence scores, and linearly projects into query tokens. We test VideoMAE (8) weights pre-trained on the EPIC-KITCHENS-100 (20) or trained from scratch. Action classes are derived by separately predicting verbs and nouns, then combining them via outer product (48, 98).
@@ F.7 Details on the Full-body Motion Generation Benchmark
For the full-body motion representation, we use redundant information to serve as input motion representation, which can usually get a more robust representation. Specifically, we concatenate both the joint locations (body, hands and eye gaze) and the joint angles (body, hands and eye gaze) as the motion representation. We obtain a 327-dim motion representation. We use our fine-grained actions as the input text, coupled with tags (e.g., $VERB$, $NOUN$, $ADV$ for adverbs) for each word. They are then transformed into word tokens using CLIP’s text encoder (92). We similarly process the middle egocentric view frame into visual features using CLIP’s image encoder (92). As baselines, we adapted two recent motion generation models, i.e., T2M-GPT (15) and MoMask (16). MoMask is the state-of-the-art model on HumanML3D (26) and T2M-GPT is another strong model working with a different mechanism. We trained models with two text prompts: verb-noun pairs (actions) and verbs only.

### v2

@@ 1 Introduction
Table 1: Action dataset comparison. #\# indicates "number of" for simplicity. The remaining columns mark following features: parametric model for motion representation (PM), structured actions (SA), markerless video recording (ML), depth recording (DR), and absolute positioning (AP). Note: EgoExo4D (21) reports 1422h by summing per camera recording time, where the total activity is 180h, yet only 88.8h annotated with MSCOCO keypoints (numbers from (31)).
@@ 3.2 Data collection procedure
To capture realistic cooking scenarios, subjects prepared a meal from reading a recipe to cleaning up, leading to significantly longer recordings than most existing action datasets (Table 1). We recruited 16 subjects (four males and twelve females, two left-handed, ages 20-46) to cook for up to five sessions in the EPFL-Smart-Kitchen (see Supp. Mat.). During each session, subjects are asked to follow one of four different recipes (omelet, pad thai, risotto, ratatouille), adapted to their preferences and requirements. Overall, we recorded and processed 29.7h of cooking experiments, corresponding to 3,207,600 frames per camera for 49 cooking sessions. All procedures were approved by the EPFL-Ethical Board. Subjects’ faces are anonymized across all videos to address privacy concerns. All subjects consented and were informed about the ethics (see Supp. Mat.).
@@ 4.4 Situated full-body motion generation benchmark
Table 4: Full-body motion generation results. Models trained and evaluated on the EPFL-Smart-Kitchen-30. FID:Fréchet Inception Distance, DIV:Diversity, $M$:Multimodality, MMd:Multimodal distance.
Pre-processing, Baselines and Metrics To achieve robust full-body motion representation, we combine joint locations and angles for the body, hands, and eye gaze, resulting in a 327-dimensional redundant motion representation. We process fine-grained action text with linguistic tags and extract egocentric visual features with CLIP’s text and image encoders (92). As baselines, we adapt three strong motion generation models, T2M-GPT (15), MARDM-SiT (93) and MoMask (16), training them with verb-noun pairs and verb-only prompts (see Supp. Mat.). Like in HumanML3D (26), we train the quantitative evaluator on EPFL-Smart-Kitchen-30 to measure the R-precision top-1 to top-3 (T1 to T3), multimodal distance (MMd), Fréchet Inception Distance (FID), Diversity (DIV), and Multimodality (MM).
Results. Our qualitative analysis (from MoMask) revealed compelling examples of successfully generated full-body motion sequences that accurately represent natural cooking behaviors (Figure 5 and Supp. Mat.). Quantitatively, we found that just like on HumanML3D(26), MoMask (16) consistently outperforms T2M-GPT (15) and MARDM-SiT (93) in different settings and metrics (Table 4.4), possibly due to the hierarchy and quantized tokenization way works better for the high dimensional properties of the whole-body representation. Furthermore, when conditioning on the egocentric view, the model was able to find some clues to minimize the global distribution and thus make the FID score lower.
@@ 5 Conclusion, future work and impact
Overall, we share multi-view, multimodal action understanding, modeling, and video question answering benchmarks to leverage the potential of multi-modality for improving action understanding and to fuel foundation models. This is particularly interesting for emerging multi-modal models (85, 86, 94).
@@ A.1 Details on devices and sensors
We installed nine Kinect Azure cameras (67) inside the kitchen and asked the subjects to wear the HoloLens 2 headset (68) to record video streams. The RGB streams were recorded at 30 FPS, and the depth streams were recorded at 10 FPS. Meanwhile, we also extracted the head position, hand poses, and eye gaze data at 30 FPS from the HoloLens 2. Apart from the video streams, we also installed six wired IMU sensors (Adafruit BNO055 sensor) on some of the appliances (fridge/cupboards) and two wireless IMU sensors (Axivity AX3 sensor) on the frequently used tools (knife/spatula). The wired IMU sensors were connected to an Arduino, which also recorded the audio signals from the Kinect Azure cameras and then sent them to those IMU sensors for synchronization.
The wireless IMU sensors were synchronized based on timestamps recorded during data collection. The wired IMU sensors were recorded at 30 FPS while the wireless IMU sensors were recorded at 100 FPS (Figure A.1 B).
@@ A.2 Synchronization and calibration of the sensors
As shown in Figure A.1 C, we synchronized the egocentric and exocentric cameras using the offset of a flashing light at the beginning of the session and use the recorded timestamps to match the captured frames. The Kinect Azure cameras can output regular audio signals; we used those signals to synchronize all the Kinect Azure cameras and the IMU sensors. We validated the calibration and synchronization by projecting the regressed 3D body poses to all the views and manually checking if the poses were aligned with the body in all the views at randomly selected timestamps.
@@ A.3 List of recipes
[…]
@@ F.1.5 Question engineering
Q6: I am currently [VERB] the [NOUN], what [BE||past,future] [NEXT/PREVIOUS] action ?
@@
Q7: I am currently [VERB], what [BE||past,future] [NEXT/PREVIOUS] action ?
@@
Q8: I am currently [ACTIVITY], what [BE||past,future] [NEXT/PREVIOUS] activity?
@@
Q11: I am currently [VERB] the [NOUN], what [BE||past,present,future] my [NEXT/CURRENT/PREVIOUS] activity ?
@@
Q12: I am currently [VERB], what [BE||past,present,future] my [NEXT/CURRENT/PREVIOUS] activity ?
@@ F.4 Details on the Action Recognition Benchmark
Data from different modalities are first sent to different modality-specific projection layers to obtain the query tokens and then sent to the transformer backbone to extract cross-modality token-wise features. Finally, features from different modalities are merged using Average Pooling to make action predictions. Specifically, for video-type data, we follow the original implementation of VideoMAE (8), select 16 frames and split into 16 $\times$ 16 patches for each frame; we introduce the egocentric view and one global exocentric view as inputs in our baseline designs. For point-type data (e.g., hand-body poses and eye gaze), we sample poses of 32 frames. For every pose, we concatenate all the pose joints and their corresponding confidences together and apply a linear projection layer to generate a query token. We tried two different variants that either use VideoMAE (8) weights pre-trained on the EPIC-KITCHENS-100 (20) or trained from scratch. We follow standard practice (48, 99) to predict verb class and noun class separately and then apply the outer product to obtain the action class.
@@ F.4.2 Data Preprocessing
We adopt VideoMAE (8) model with the VIT-L backbone (100) to build our multimodal baselines. Data from different modalities are projected into query tokens, processed by a transformer for feature extraction, and merged via Average Pooling for action prediction. Video data uses VideoMAE (8) with 16 frames, incorporating egocentric and global exocentric views. Point-type data (e.g., hand-body poses, eye gaze) samples 32-frame poses, concatenates joint positions with confidence scores, and linearly projects into query tokens. We test VideoMAE (8) weights pre-trained on the EPIC-KITCHENS-100 (20) or trained from scratch. Action classes are derived by separately predicting verbs and nouns, then combining them via outer product (48, 99).
@@ F.7 Details on the Full-body Motion Generation Benchmark
For the full-body motion representation, we use redundant information to serve as input motion representation, which can usually get a more robust representation. Specifically, we concatenate both the joint locations (body, hands and eye gaze) and the joint angles (body, hands and eye gaze) as the motion representation. We obtain a 327-dim motion representation. We use our fine-grained actions as the input text, coupled with tags (e.g., $VERB$, NOUNNOUN, ADVADV for adverbs) for each word. They are then transformed into word tokens using CLIP’s text encoder (92). We similarly process the middle egocentric view frame into visual features using CLIP’s image encoder (92). As baselines, we adapted two recent motion generation models, i.e., T2M-GPT (15) and MoMask (16). MoMask is the state-of-the-art model on HumanML3D (26) and T2M-GPT is another strong model working with a different mechanism. We trained models with two text prompts: verb-noun pairs (actions) and verbs only.